OpenStudy (anonymous):

Angle C is an inscribed angle of circle P. Angle C measures (20x – 5)° and arc AB measures (30x + 30)°. Find the measure of arc AB.

OpenStudy (jdoe0001):

 \(\bf \textit{incribed angle}=\cfrac{\textit{intercepted arc}}{2}\implies 2\textit{incribed angle}=\textit{intercepted arc}\)

OpenStudy (anonymous):

so I put 30x+30 over 2 ? then solve it from there

OpenStudy (jdoe0001):

yes, you can do it that way

OpenStudy (anonymous):

I keep getting 18 and I know that isn't the answer im very confused

OpenStudy (jdoe0001):

\(\bf \textit{incribed angle}=\cfrac{\textit{intercepted arc}}{2}\implies 2\textit{incribed angle}=\textit{intercepted arc} \\ \quad \\ 20x-5=\cfrac{30x+30}{2}\implies 40x-10=30x+30\implies 10x=40 \\ \quad \\ x=4\)

OpenStudy (jdoe0001):

does that give you 18?

OpenStudy (anonymous):

I kind of get it now but when I do it myself its doesn't come out the way you do

OpenStudy (jdoe0001):

2(20x-5) distributing the 2, will give you 40x-10

OpenStudy (anonymous):

oh ok I get it now thank you so much I appreciate your help

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

Point C is the center of the circle. If angle ACB measures 4 x plus 12 degrees and AB measures x plus 13 degrees, find x. http://curriculum.kcdistancelearning.com/courses/ALG2x-CR-A09/b/assessments/G-GeometryUnitExam/A2CR_Unit7_Exam_LT_118q3.gif